A Note on Certain Bibasic q-Appell and Lauricella Series

Bhaskar Srivasatva, Lucknow, India

Keywords: q-Bibasic Hypergeometric series, q-Appell and q-Lauricella series AMS Subject classification: 33D

Abstract: On this note we have given certain transformations of q-Appell and q-Lauricella series on two bases. It is interesting to note that the q-Appell series on two bases can be reduced to an expression with only one base. Limiting cases of q-Appell series are closely connected with Ramanujan’s G(a, λ) function and so has been expresed in terms of continued fractions.